### v2

# villa-X: Enhancing Latent Action Modeling in Vision-Language-Action Models

[Uncaptioned image]Äpgjy
villa-X: Enhancing Latent Action Modeling in Vision-Language-Action Models
Xiaoyu Chen Hangxing Wei Pushi Zhang Chuheng Zhang Kaixin Wang Yanjiang Guo Tsinghua University Rushuai Yang Yucen Wang Xinquan Xiao Li Zhao Jianyu Chen Tsinghua University Jiang Bian Microsoft Research
Visual-Language-Action (VLA) models have emerged as a popular paradigm for learning robot manipulation policies that can follow language instructions and generalize to novel scenarios. Recent work has begun to explore the incorporation of latent actions, an abstract representation of visual change between two frames, into VLA pre-training. In this paper, we introduce villa-X, a novel Visual-Language-Latent-Action (ViLLA) framework that advances latent action modeling for learning generalizable robot manipulation policies. Our approach improves both how latent actions are learned and how they are incorporated into VLA pre-training. Together, these contributions enable villa-X to achieve superior performance across simulated environments including SIMPLER and LIBERO, as well as on two real-world robot setups including gripper and dexterous hand manipulation. We believe the ViLLA paradigm holds significant promise, and that our villa-X provides a strong foundation for future research.
Keywords: Latent Action, Vision-Language-Action Model code: https://github.com/microsoft/villa-x site: https://aka.ms/villa-x
1 Introduction
Latent action learning has emerged as a promising approach for the pretraining of vision-language-action (VLA) models [5, 31, 52, 11, 30, 68, 35, 13], enabling learning from both robot data and human video data [67, 11, 52, 1]. At the core of these methods is a Latent Action Model (LAM), which compresses visual changes between successive frames into latent tokens, referred to as latent actions. These latent actions are expected to capture motion semantics. These tokens function as pseudo-action labels for imitation learning and enriching robot policy training with abundant, action-free data.
Yet two questions remain: how to learn latent actions and how to incorporate latent actions into VLA pre-training? In this paper, we introduce villa-X, a novel Visual-Language-Latent-Action (ViLLA) framework that advances both key aspects of latent action modeling. For the latent action learning component, a key limitation is the absence of direct alignment between visual transitions and underlying robot states and actions. Although robot data with low-level states and actions are often available alongside action-free videos during pretraining, current methods choose to ignore these signals and focus solely on the visual part. As a result, latent actions can remain ungrounded in the robot’s physical dynamics, resulting in a weak correspondence between latent action and robot behavior. To address this issue, we incorporate a proprio Forward Dynamics Model (FDM) module into the LAM as an auxiliary decoder. This module predicts future robot proprioceptive states and actions, encouraging the learned latent actions to better reflect the agent’s behavior—not just the observed visual changes. Moreover, this design improves the interpretability of the learned latent actions and enables them to be more easily translated into executable robot actions. As a result, latent actions become a robust intermediary between vision-based representations and low-level controls. To exploit this structure, we propose to model the latent action and robot action distribution jointly through a joint diffusion process, and conditioning robot action generation on latent actions generation through attention. Compared to existing methods on latent action pre-training, our approach allows for more effective and structured information transfer from latent actions to robot actions.
We conduct comprehensive evaluations of villa-X on two simulated environments, SIMPLER and LIBERO, as well as two real-world setups, including various robot platforms with gripper and dexterous hand manipulation. Our results demonstrate that villa-X achieves superior performance, laying a robust foundation for future research in this domain.
Our contributions are listed as follows:
  • We improve latent action learning by introducing an extra proprio FDM, which aligns latent tokens with underlying robot states and actions and grounds them in physical dynamics.
  • We propose to jointly learn a latent action expert and a robot action expert through joint diffusion in the policy model, conditioning robot action prediction on latent actions to fully exploit their potential.
  • Our method demonstrates superior performance on simulated environments as well as on real-world robotic tasks. The latent action expert can effectively plan into future with both visual and proprio state planning.
Figure 1: (a) High-level architecture of our villa-X method. (b) Detailed structure of the Latent Action Model (LAM) component within villa-X.
2 Related Work
Visual-Language-Action Models
Vision-Language-Action (VLA) models [5, 31, 52, 11, 30, 68, 35, 13] leverage pre-trained vision-language models (VLMs) to generate robot actions using visual and language cues. They either directly repurpose VLMs for action prediction [11, 67, 35, 30] or use action experts to map VLM outputs to robot actions [52, 5, 31]. While training on large-scale datasets like Open X-Embodiment [13] enhance the generalization ability of VLAs, cross-embodiment generalization remains challenging due to diverse robot setups and configurations. Utilizing unlabeled trajectory data with pseudo-labels such as latent actions [67, 11], language sub-goals [49], or visual sub-goals [68] supports overcoming these challenges. Our method adopts the latent action approach and improves both the modeling of latent actions and their integration into VLA pretraining.
Modeling Latent Actions for VLA Pretraining
Recent exploration into latent actions began with LAPO [58] and Genie [7], which primarily focused on the video game domain. Dynamo [15] adopted a similar architecture, using inverse and forward dynamics models to shape state representations. LAOM [51] propose to use supervision to learn better latent actions in the presense of distractors on Mujoco environments.
For robotic learning, methods have started to incorporate latent actions into VLA pretraining [67, 10, 11, 52, 1, 8, 66, 38]. LAPA [67] proposes to learn from videos, and trains its latent actions and Vision-Language Model (VLM) using either human or robot video data. Concurrently, IGOR [10] learns latent actions from a mixture of human and robot videos, marking the first demonstration of successful latent action transfer between humans and robots in a unified action space for embodied AI. Moto-GPT [11] co-fine-tunes both latent and robot action labels. GR00T [52] treats latent actions as a distinct embodiment, while Go-1 [1] generates robot actions conditioned on discrete latent tokens. UniVLA [8] proposes a two-stage training pipeline to learn task-centric latent actions. More recent works like [66, 38] explore the continuous latent actions. By contrast, our approach jointly models latent and robot actions through a joint diffusion process, conditioning robot action generation on latent actions for more effective and structured information transfer. Our method offers better integration compared to LAPA [67] and GR00T [52], incorporates immediate visual context unlike Moto-GPT [11], and avoids inconsistencies caused by teacher forcing seen in Go-1 [1], achieving robust test-time reasoning.
3 Method
Figure 2: Visualization of image pairs with similar latent actions.
At a high level, our villa-X consists of two main components, as illustrated in Figure 1a:
  1. (i)
    a LAM (Latent Action Model) module that infers latent actions from a pair of observations;
  2. (ii)
    a ACT (Actor) module that jointly models latent actions and robot actions given an initial visual observation and a textual task instruction.
We describe each component in detail in the following subsections. Learning in villa-X is carried out in three stages: (1) pretraining the LAM module, (2) pretraining the ACT module, and (3) finetuning ACT on target-embodiment data.
3.1 Latent Action Model
Modeling latent actions enables the use of abundant action-free video demonstration data and also holds the promise of improving generalization between robot and human motions [67, 11]. Building on these prior works, our LAM (depicted in Figure 1b) learns an inverse dynamics model (IDM) to extract latent actions $z_{t}$ from two adjacent video frames, $o_{t}$ and $o_{t+K}$. The interval $K$ is a configurable window size, chosen to balance perceptibility and generalizability: large enough to capture meaningful motion but small enough to maintain transferability.
Previous works [67, 11] train IDM by jointly optimizing a forward dynamics model (FDM) to predict future frames. Given the current frame $o_{t}$ and a latent action $z_{t}$, the model minimizes the discrepancy between the predicted frame $\hat{o}_{t+K}$ and the ground truth $o_{t+K}$. Although robot data with low-level states and actions are often available alongside action-free video data, these approaches completely discards such information, focusing solely on visual changes. Since our ultimate objective is to use latent actions to facilitate real robots policy learning, bridging vision-language prompts and low-level controls would be beneficial. To this end, we propose augmenting LAM pretraining with proprioceptive supervision. Specifically, since $z_{t}$ represents a higher-level action abstraction, we introduce an additional proprio FDM module that predicts future robot states and actions $K$ steps ahead given the current robot state $q_{t}$ and the latent action $z_{t}$. This alignment of high-level abstractions and physical robot dynamics enhances the suitability of latent action for robot policy learning.
However, because robot trajectories are often embodiment specific, conditioning solely on $q_{t}$ and $z_{t}$ may embed platform-dependent features in the latent action, hindering cross-embodiment transfer. To alleviate this issue, we provide the proprio FDM with an additional embodiment context input (e.g., dataset ID, control frequency), isolating transferable action representations from embodiment particulars.
[Uncaptioned image] Implementation
In practice, instead of predicting a single latent action between two frames (as Figure 1b depicts), LAM receives a sequence of $T_{\text{LAM}}$ observation frames as input and outputs $T_{\text{LAM}}-1$ latent actions between consecutive frames. The inverse dynamics model first encodes the frame sequence with a Spatial–Temporal Transformer [65], then applies value quantization to produce $T_{\text{LAM}}-1$ latent actions $z\in\mathbb{R}^{N_{z}\times D_{d}}$. The FDM is implemented using a Vision Transformer (ViT), while the proprio FDM is a two-layer MLP with dual output heads. The embodiment context input is embedded using learnable embeddings and concatenated with the robot state before being fed into the proprio FDM. For further implementation details and configuration specifics, please refer to the appendix.
🖹 Training and Inference
We train LAM on diverse data sources, including robot trajectories and human egocentric videos. For data lacking low-level state and action labels, we omit the proprio FDM branch and optimize only the FDM reconstruction loss. By leveraging this diverse data, our goal is to learn a generalizable intermediate representation for downstream policy learning. Once trained, only the IDM is required to extract latent actions from a pair of frames at inference, the FDM and proprioceptive FDM serve primarily as diagnostic visualization tools. We use a mixture of datasets containing both human video and robot action data including OpenX [13], Ego4D [22], etc. For more details, please refer to Appendix A.1.
3.2 Actor Module
Figure 3: Structure of the ACT module.
In the same vein as common VLA models [30, 31, 5, 67, 52], our ACT module builds upon a pre-trained vision-language model (VLM) and is trained to model a generalizable robot control policy that outputs actions given a textual task instruction and an initial observation. With the pre-trained LAM, we are able to incorporate latent actions as an additional intermediate representation in our framework, rather than directly mapping vision and language inputs to low-level robot actions. The latent actions serve as a bridging language that connects task prompts to robot action planning.
Figure 3 provides an illustration of our ACT module, which consists of three components: VLM, ACT-latent, and ACT-robot. Given a textual task prompt $l$ and a visual observation $o_{t}$, the VLM first extracts visual and language features. Then, the ACT-latent component, trained with a diffusion objective, models a sequence of $n$ latent actions $z_{t},z_{t+K},\ldots,z_{t+(n-1)K}$, conditioned on the features from the VLM. Subsequently, the ACT-robot component, also implemented as a diffusion model, predicts a sequence of $m$ low-level robot actions, conditioned on the same visual and language features as well as the latent actions. Conditioning across components is achieved using uni-directional attention over intermediate features, as indicated by the dashed line in Figure 3. Additionally, ACT-robot can incorporate wrist camera input when available. Unlike prior work [5], which feeds wrist camera data into the VLM from the start, our design is more computationally efficient. Another consideration is that we view the wrist camera input as more embodiment-specific and less generalizable, which is why we incorporate it into ACT-robot rather than the VLM. The first two components, VLM and ACT-latent, are designed to handle the more generalizable aspects of the policy.
One might ask whether the length of the robot action sequence must match that of the latent action sequence, i.e., whether $m=nK$. In our design, this is not required. We allow different lengths for latent and robot action sequences, as long as they share the same starting time step $t$. For example, we may model 10 latent actions (corresponding to 20 robot actions if $K=2$) but choose to predict only 4 robot actions (effectively relying on the first 2 latent actions). This flexibility enables us to maximize the benefits of modeling latent actions.
We highlight three key elements of our design for incorporating latent actions:
  • Latent actions as a mid-level bridge. Previous works, including LAPA [67] and GR00T [52], treat latent actions similarly to robot actions and do not explicitly model the hierarchical structure between them. For example, GR00T simply treats latent actions as an additional action type and trains them alongside robot actions within a shared diffusion transformer. In contrast, our model treats latent actions as a distinct mid-level representation that bridges high-level vision and language prompts with low-level robot actions.
  • Explicit transfer from latent to robot actions. Unlike LAPA [67] and GR00T [52], where the transfer from latent actions to robot actions occurs only implicitly through pre-trained weight initialization, our model enables an explicit connection. The robot action diffusion process is directly conditioned on the latent action diffusion process, allowing more effective and structured information transfer from latent actions to robot actions.
  • Modeling latent actions sequences. While prior works (LAPA [67] and GO-1 [1] ) incorporate latent actions, they only model one-step predictions rather than planning over longer horizons at the latent action level. In contrast, our approach models a sequence of future latent actions, enabling structured planning at both the latent and robot action levels.
Implementation  We use a pre-trained PaliGemma model [3] as our VLM. Both ACT-latent and ACT-robot adopt a transformer architecture following [69], and are trained using flow matching [39, 42]. Wrist camera images are tokenized following the approach used in HPT [63]. Similar to the proprio FDM in LAM, we incorporate embodiment context embeddings in ACT-robot to account for embodiment-specific information. Inspired by Moto [11] and RDT [43], we apply attention masking during training: with 50% probability, we completely disable attention from latent actions to robot actions; otherwise, we randomly mask out attention from 50% of the latent actions. Further implementation details are provided in the appendix.
🖹 Training, Finetuning and Inference During training, we update all components of the model, including both the pre-trained and randomly initialized parts. The model is trained jointly using losses on both latent actions and robot actions, leveraging a mixture of datasets containing both human video and robot action data including OpenX [13], Ego4D [22], etc. Finetuning follows the same procedure as training, with the exception that, for unseen embodiments, we initialize a new context embedding. For robot action inference, we perform denoising steps simultaneously in both ACT-Latent and ACT-Robot. For more details, please refer to Appendix A.2.
4 Experiments
In this section, we aim to answer the following questions through experiments:
  • Does our improved LAM learn higher-quality latent actions?
  • Can ACT-latent successfully plan future motions?
  • Can the actor module effectively leverage the pre-trained latent actions?
  • How does villa-X compare to existing VLA baselines in both simulated benchmarks and real-world robot tasks?
4.1 Does our improved LAM learn higher-quality latent actions?
In this subsection, we evaluate whether our improved latent action modeling enhances the quality of the learned latent actions. The key component of our LAM is the incorporation of the proprio FDM module. To assess its impact, we compare our model (denoted w/pp) to a variant without the proprio FDM module (denoted wo/pp).
Probing  First, a core expectation for latent actions is that they should carry information useful for predicting low-level robot actions. To test this, we conduct a probing experiment. Specifically, after training the latent action models, we freeze them and train a simple 3-layer MLP to predict the corresponding robot actions for each latent action. Probing is conducted on the LIBERO dataset [40], which is not used for training latent action models. We train the MLP on the training split of LIBERO and evaluate it using the L1 loss on the validation split.
Figure 4: Probing experiment results.
We are interested in how closely the predicted actions match the ground truth. In LIBERO, the robot action space has 8 dimensions: 3 for position, 4 for rotation, and 1 for the gripper. Rather than averaging the error across dimensions, we focus on the maximum L1 error across all action dimensions, as we want to avoid large deviations in any single aspect of the action. For each model variant (w/pp and wo/pp), we compute the number of validation samples whose maximum L1 error falls below a threshold. By sweeping this threshold, we count how many samples fall within each error bin. A better model should yield more samples with low errors.
For each error bin, we compute the difference in the number of samples between the w/pp and wo/pp variants and present the results in Figure 4. The w/pp variant produces more samples with smaller errors, while the wo/pp variant has more samples in the high-error bins. This demonstrates the effectiveness of the proprio FDM module in capturing information from the robot actions.
Table 1: Evaluation results on SIMPLER for different variants of our villa-X (top group) and alternative approaches for incorporating latent actions (bottom group). ”Ours” refers to the w/pp described in the main text.
Method Google robot WidowX robot Pick Move Drawer Place Avg. Carrot Eggplant Spoon Cube Avg. Ours 81.7 55.4 38.4 4.2 44.9 24.2 71.7 48.3 19.2 40.8 wo/pp 77.0 52.7 42.6 2.8 43.8 22.5 57.5 43.3 5.9 32.3 wo/LAM 42.1 24.6 38.4 0.0 26.3 25.8 60.8 36.7 9.2 33.1 LAPA-style 64.7 28.8 38.0 5.6 34.2 0.8 0.0 2.5 0.8 1.0 Go-1-style 29.0 38.0 31.3 4.6 25.7 5.8 50.8 1.7 1.0 14.8
Policy Pre-training  Next, we compare how the latent actions generated by different variants of LAM (w/pp and wo/pp) influence policy pre-training. Unlike the main experiments, we pretrain models in this section on a mixture of 10% Fractal [6] data, 10% Bridge V2 [18] data, and 100% Something-Something V2 [20] data, to reduce computation cost while remaining a setting where limited robot data are available for training the VLA model. The resulting policies are evaluated in the SIMPLER environment [32], a simulation benchmark explicitly designed to mitigate the gap between simulated and real‐world robotic environments. It comprises two platforms, namely Google robot and WidowX robot, each with four manipulation tasks. We evaluate our method on all the 8 tasks on the visual matching setting. The results are summarized in Table 1. We observe that w/pp clearly outperforms wo/pp, demonstrating the effectiveness of incorporating the proprio FDM module. Additionally, we include a baseline that does not use latent actions (denoted wo/LAM) and is trained solely to predict robot actions. The performance of wo/LAM is significantly worse, indicating that pre-training with latent actions is essential.
LAM visualization
Figure 2 visualizes image pairs sharing the same latent action, demonstrating that these pairs correspond to similar underlying robot behaviors. To further demonstrate the transfer ability of our LAM, we extract latent actions from arbitrary video sequences, map them to robot actions using the proprio FDM, and execute the resulting robot action in SIMPLER simulator. The simulated motions closely reproduce the original demonstrations, indicating that latent actions are both aligned with and grounded in the robot’s actions. For more examples, please check Appendix C.1.
Figure 5: Transfer of human demonstrations to new scenes via latent actions and a world model. (a) moving right. (b) moving downwards. Please visit our website for more demos.
Figure 6: Transfer of video demonstrations into robot action executions via latent actions and a proprio FDM. Please visit our website for more demos.
LAM transfer consistency
We provide two visualizations to further evaluate the transfer consistency of the learned latent actions. Specifically, we first use the trained LAM to extract a sequence of latent actions from a source video. These latent actions are then applied to different initial images, allowing a world model to generate subsequent frames conditioned on these actions. As shown in Figure 5, the transfer rollouts (highlighted in blue) successfully identify the movable parts and demonstrate behaviors consistent with the original video.
We also evaluate the alignment between latent and robot actions. To do this, we use the proprio FDM to decode latent actions into executable robot actions. These actions are then executed in the SIMPLER simulator, with the resulting rollouts visualized in Figure 6. In each example, the top row shows the original video demonstration from which latent actions were extracted, while the bottom row displays the corresponding simulated execution. The simulated motions closely match the original demonstrations, confirming that the latent actions learned by villa-X are well-aligned with and grounded in the robot actions.
4.2 Can ACT-latent successfully plan future motions?
Figure 7: Transfer of video demonstrations into robot action executions via latent actions and a proprio FDM. Please visit our website for more demos.
In this experiment, we demostrate the motion planning capabilities of ACT-latent by visualizing its planned actions with a world model. For a given initial image and language instruction, ACT-latent generates a sequence of latent actions, and the world model takes the inital frame and latent actions as inputs to render the latent actions into planned future videos. Figure 7 shows results for both in-distribution and out-of-distribution samples, where the in-distribution samples are randomly sampled from the validation set but from the same datasets as used in training, and the out-of-distribution samples are from new real-world scenarios. The results demonstrate that ACT-latent successfully follows the language instructions to solve the tasks, while accurately identifying target objects and generating latent actions that successfully follow the instructions. ACT-latent also successfully identifies the concepts in the emoji, which rarely appears in robot datasets, suggesting that villa-X keeps the general vision-language capabilities in the initial VLM model after pre-training.
4.3 Can the actor module effectively leverage the pre-trained latent actions?
Given high-quality latent actions produced by the pre-trained LAM, we investigate whether our design can effectively leverage them to pre-train robot control policies. We compare our approach against two recent methods that also utilize latent actions, albeit in different ways: LAPA [67] and GO-1 [1].
To isolate the effect of how latent actions are incorporated, we implement LAPA-style and GO-1-style models based on our architecture for a fair comparison. For the LAPA-style model, we follow a two-stage pre-training protocol: we first train the VLM to predict latent actions, then replace the latent action prediction head with a robot action prediction head and continue training on data with robot action labels. For the GO-1-style model, we implement a separate latent planner that autoregressively predicts latent actions. The robot action prediction component remains largely unchanged as in our main design.
Following the experiment setup in the previous subsection, we train all models on the same dataset mixture and then evaluate the resulting policies in the SIMPLER environment [32]. The results are shown in Table 1. Compared to other two approaches, our method achieves significantly higher performance, validating the effectiveness of our design for incorporating latent actions into VLA pre-training.
4.4 Evaluating villa-X in Simulation
4.4.1 SIMPLER Benchmark
Table 2: Comparison on SIMPLER of villa-X and existing methods. Methods marked with * are evaluated directly after pretraining, whereas other methods are evaluated after post-training on corresponding dataset.
Method Google Robot WidowX Robot Pick Move Drawer Place Avg. Carrot Eggplant Spoon Cube Avg.. RT-1-X * 56.7 31.7 59.7 21.3 42.4 4.2 0.0 0.0 0.0 1.1 Octo-base * 17.0 4.2 22.7 0.0 11.0 8.3 43.1 12.5 0.0 16.0 OpenVLA * 16.3 46.2 35.6 0.0 24.5 0.0 4.1 0.0 0.0 1.0 RoboVLMs * 72.7 66.3 26.8 36.1 50.5 25.0 0.0 20.8 8.3 13.5 RoboVLMs 77.3 61.7 43.5 24.1 51.7 20.8 79.2 45.8 4.2 37.5 GR00T-N1.5 69.3 68.7 35.8 4.0 44.5 54.3 61.3 75.3 57.0 62.0 MoTo 74.0 60.4 43.1 N/A N/A N/A N/A N/A N/A N/A LAPA N/A N/A N/A N/A N/A 45.8 58.3 70.8 54.2 57.3 Ours w/o latent 56.3 25.8 27.3 13.9 30.8 31.3 74.6 61.7 28.3 49.0 Ours 98.7 75.0 59.3 5.6 59.6 46.3 64.6 77.9 61.3 62.5
Baselines and Experimental Setup
We use the SIMPLER benchmark as described above. In this section, we compare against two categories of prior work:
  • Vision-Language-Action (VLA) models: RT-1-X [13], Octo-base [54], OpenVLA [30], RoboVLMs [34], which learn policies solely from mixed robot datasets.
  • Joint policy learning and world modeling method: GR00T-N1.5  [53], which aligns the model with target future embeddings.
  • Latent-Action based methods: MoTo [11] and LAPA [67], which additionally exploit unlabelled videos by inferring latent actions.
Except where noted (*), all models follow a two‐stage pretraining–finetuning protocol, including a general pretraining phase on large‐scale mixed data, followed by finetuning on a dataset of specific embodiment. We also include an ablation (villa-X w/o latent) that removes our latent‐action expert while keeping all other components unchanged. Reported baseline scores are taken directly from the original publications, while missing entries are marked as N/A.
Experimental Results
Table 2 summarizes the success rates on both platforms. Our full model achieves the highest score on average success rate on the Google robot (59.659.6%) and the WidowX robot (62.562.5%). This improvement over VLA methods, which cannot exploit unlabelled video, demonstrates the benefit of our incorporating human videos into policy learning. Moreover, our approach outperforms other latent‐action methods, indicating that our specific mechanism for leveraging video data is more effective. Finally, the gap between our full model and the “villa-X w/o latent” ablation confirms that the latent‐action expert is essential for achieving these gains.
4.4.2 LIBERO Benchmark
The LIBERO benchmark [40] evaluates knowledge transfer in multitask and lifelong robot learning problems for robotic manipulation, consisting of four task suites: LIBERO-Spatial evaluates the model’s performance under novel layouts with the same task and object types, LIBERO-Goal evaluates the model’s performance under novel tasks with the same object types and layouts, LIBERO-Object evaluates the model’s performance under novel object types with the same tasks and layouts, LIBERO-Long evaluates the model’s performance under diverse set of objects, layouts and backgrounds. Each task suite contains 10 tasks with 50 human demonstrations per task for fine-tuning.
Baselines and Experimental Setup
We compare with the following existing models: Diffusion Policy [12] trained from scratch, Octo [54], and OpenVLA [30]. All models follow a two-stage pretraining-finetuning protocol. We finetune villa-X and villa-X w/o latent on the demonstration data of the each task suite separately, and test on the LIBERO simulator for 10 tasks and 20 trials per task on each task suite.
Experimental Results
Table 3 summarizes the success rates on each task suite of LIBERO. Our model achieves better performance than existing methods in all the four task suites. Also, our model with latent action achieves higher performance on 3 of the 4 task suites and average performance, confirming that the proposed latent action expert improves the manipulation performance.
Table 3: Evaluation on 4 LIBERO task suites of villa-X and existing methods.
Method Spatial Object Goal Long Average Diffusion Policy [12] 78.3 92.5 68.3 50.5 72.4 Octo-base [54] 78.9 85.7 84.6 51.1 75.1 OpenVLA [30] 84.7 88.4 79.2 53.7 76.5 Ours w/o latent 86.0 86.5 85.0 70.0 81.9 Ours 97.5 97.0 91.5 74.5 90.1
4.5 Evaluating villa-X on Real-world Robots
Figure 8: Real-world robot evaluation platforms: (top) Realman robot arm platform with a gripper and (bottom) Xarm robot arm with Xhand dexterous hand. Platform setups are shown on the left, with corresponding evaluation tasks on the right. Please visit our website for more videos.
Table 4: Evaluation on Xarm robot arm of villa-X and existing methods.
Method Pick & Place Stack Cube Place Cup Upright Pour Water Flick Ball seen unseen seen unseen seen unseen seen unseen seen unseen GR-1 56 40 15 5 0 0 0 0 40 10 GR00T 44 28 20 0 20 0 0 0 30 0 Ours w/o latent 72 60 70 40 40 30 40 10 50 30 Ours 84 68 75 50 60 30 60 30 50 40
Table 5: Evaluation on Realman robot arm of villa-X and existing methods.
Method Pick in Pick out Push Stack Unstack Change block color Change table cover GR00T 30 70 10 10 60 50 30 Ours w/o latent 40 80 30 60 70 40 30 Ours 30 100 50 50 100 60 60
To further assess generalization, we evaluate our approach on two real‑world robotic platforms: a Realman robot arm with a gripper, and an Xarm manipulator equipped with a 12‑DOF Xhand dexterous hand. See Figure 8 for details.
Realman robot arm with gripper
On the robot-arm manipulation platform, we use a 6-DoF Realman RM 75 robot arm and a 1-DoF Inspire gripper. We fine-tune and evaluate our policy on “Pick-in” (pick the block into a bowl), “Pick-out” (pick the block out of a bowl), “Stack” (stack the block onto another block), “Unstack” (unstack the block from another block), and “Push” (push the block to a given location) tasks. For fine-tuning, we collect a dataset of 375 trajectories through teleoperation (with 75 trajectories for each task), where the object layout and table setup are fixed and only the object locations are dynamic.
We conduct two sets of evaluation: In task evaluation, we remain the table setup the same as data collection; in generalization evaluation, we change the color of the block and table cover. We evaluate the policy with 10 trials per task with each trial conducted with different object location, where all the experiment settings are the same for the evaluation of different policies (including the object location and lightening condition). The results are shown in Table 5, demonstrating that villa-X outperforms existing baselines in both task and generalization evaluation sets.
Xarm robot arm with Xhand dexterous hand
On the dexterous‑hand platform, we use the Xhand, a 12‑DoF dexterous hand with five flexible fingers, mounted on a 7-DoF Xarm robot arm. Fine‑tuning is performed on the Xhand Dataset [26], which comprises 4,000 trajectories spanning 13 task categories. Since no dexterous-hand data were used during pretraining, this evaluation can test embodiment transfer ability. We select five representative tasks—pick‑and‑place, cube stacking, cup upright placement, water pouring and ball flicking. The results are summarized in Table 4 for (i) seen tasks, where objects are randomly replaced or additional distractors are added, and (ii) unseen tasks, which use unseen objects or backgrounds. The performances are evaluated under 50 runs for pick and place, 20 runs for stack cube and 10 runs for others. Table 4 demonstrates that our method outperforms existing baselines.
5 Conclusion, Limitations, and Future Works
In this paper, we presented villa-X, a novel Visual-Language-Latent-Action (ViLLA) framework that improves both the learning of latent actions and their incorporation into VLA pre-training. Our experiments demonstrate that our enhanced Latent Action Model learns higher-quality latent actions, and our improved policy model more effectively leverages these learned actions. Overall, our method exhibits superior performance in both simulated environments and real-world robotic tasks.
One limitation is that the proposed latent expert, although effective at future planning through both visual and proprioceptive state planning, is not fully explored in this work. For example, future research could learn a critic with prior knowledge from foundation vision-language models, allowing multiple samples from the latent expert and rejecting planned trajectories that do not follow the language instruction. We leave this aspect as future work to further improve the capability of the ViLLA framework.
Appendix A Model Design
A.1 Latent Action Model
There are four modules in our latent action model. A ST-Transformer based inverse dynamic model (IDM), a vector quantization module, an image reconstruction forward dynamic model (FDM) and a robot proprio FDM. The IDM takes an 8 × 224 × 224 video clip as input, following a Patch Embedding with a patch size of 14 and 12 ST-blocks, where each block has a hidden dimension of 768 and 32 attention heads. The codebook size of vector quantization is set to 32. Image reconstruction FDM is a 12-layer ViT-base network, and robot proprio FDM is a 2-layer MLP with a hidden size of 768. Our model is trained with a learning rate of 1.5e-4, a batch size of 512, and a 2000-step linear warmup. We use the same loss weight for the image FDM and proprio FDM. The pretraining takes 4 days on 128 NVIDIA A100 GPUs.
A.2 Actor Module
Our VLA model comprises three components. First, the vision–language encoder is based on PaliGemma [4], a 3B-parameter VLM pretrained with 224 × 224 images and 128-token text inputs. Second and third, the latent-action expert and the robot-action expert are each implemented as 18-layer Transformer networks, mirroring PaliGemma’s design, with a hidden dimension of 1,024 and 8 attention heads. For the latent action sequence, we select a sequence length of $N=6$, and for the robot actions, we select a sequence length of $M=4$.
We extend our policy head with a variant of HPT [62], assigning each embodiment its own pair of state‐ and action‐projection layers while sharing all other parameters. Visual features from the wrist camera are extracted by a pretrained ResNet-18 [24] and fused into the main model via a shared cross-attention head that maps the ResNet features into $16$ tokens. During training, wrist-view inputs are randomly masked 50% of the time. We also observed that the latent‐action representation can be overly exploited by the robot‐action expert, so we regularize this with two complementary dropout schemes. First, we add a 50% attention-weight dropout on the latent‐action stream. For the remaining tokens, we randomly mask 50% latent action tokens. This combined masking strategy encourages the model to learn robust, generalizable policy that will balance the predicted latent actions as well as the input image and instruction. Each expert contains approximately 300 M parameters and is trained from scratch. We train all components jointly using a learning rate of 5e55e-5 with a $200$-step linear warmup. We clip gradients to a maximum norm of 1.0 to ensure stable optimization. The pretraining takes 4 days on 64 NVIDIA A100 GPUs.
Appendix B Dataset
B.1 Data Mixture
We curated a data mixture by combining both robot data and action-free human videos for our pretraining phase. For robot data, we draw primarily from OpenX [13] mixture and AgiBot [1]. For OpenX dataset, our base data mixture is created primarily based on [30, 54]. In total, we use 223.5M trajectories with 1.6M frames of robot data. For human videos, we use a mixture of Ego4D [22], EgoPAT3D [37], EGTEA Gaze+ [36], EPIC-KITCHENS [16], HO-Cap [61], HOI4D [44], HoloAssist [64], RH20T [19], Something Something V2 [21]. Altogether, this yields 3.6M clips of human videos. During LAM pretraining, we exclusively utilize the primary third-person camera view. For policy pretraining, we optionally incorporate the wrist-mounted view (when available), applying a 5050% dropout. A full breakdown of our data mixture is listed in Table 6.
B.2 Data Preprocessing
For data cleaning, we adopt EgoHOD [55], a curated subset of Ego4D [22], and further filter the videos based on visual quality to ensure high-quality inputs for training. For both robot data and human videos, we apply random adjustments to brightness, contrast, saturation, and hue as data augmentation. In the case of robot data, we represent both proprioceptive states and actions using euler angles.
Dataset Mix Ratio (%)
RT-1 Robot Action [6] 9.70
AgiBot World Beta [1] 20.0
Kuka [28] 1.97
Bridge [60, 18] 5.47
Taco Play [57, 47] 0.76
Jaco Play [17] 0.12
Berkely Autolab UR5 [9] 0.31
Language Table [46] 0.11
Stanford Hydra Dataset [2] 1.61
NYU Franka Play Dataset [14] 0.22
Furniture Bench Dataset [25] 0.63
Austin Sailor Dataset [50] 0.57
Austin Sirius Dataset [41] 0.45
BC-Z [27] 3.47
DLR EDAN Shared Control [56] 0.01
CMU Stretch [48] 0.04
FMB Dataset [45] 0.73
DobbE [59] 0.37
DROID [29] 3.46
Ego4D [23, 55] 21.46
EgoPAT3D [37] 0.94
EGTEA Gaze+ [36] 0.89
EPIC-KITCHENS [16] 6.95
HO-Cap [61] 0.63
HOI4D [44] 1.99
HoloAssist [64] 4.77
RH20T [19] 5.56
Something-Something V2 [20] 6.82
Table 6: Our training data mixture used during the pretraining phase.
Appendix C LAM visualization results
C.1 Image Pairs with Similar Latent Actions
Figure 9: Visualization of image pairs with similar latent actions.
Figure 9 visualizes additional image pairs sharing the same latent action, demonstrating that these pairs correspond to similar underlying robot behaviors.
The results demonstrate that similar latent actions represent the similar robot behaviors and low-level actions, in regardless of which embodiment (including human and different robots) is executing such action. This results support that villa-X learns cross-embodiment prior knowledge for manipulations with latent actions.
Appendix D Simulation Evaluation Details
D.1 SIMPLER Benchmark
We evaluate on all eight SIMPLER [33] tasks in the visual matching setting, which include two robot platforms: Google Robot and WidowX.
For Google Robot, the tasks are: (1) pick coke can (including horizontal, vertical and standing can configurations); (2) move an object near a target object; (3) open / close top, middle or bottom drawer; and (4) place apple in a closed drawer, which includes two subtasks: first open top drawer, and then place the apple into the top drawer. On the widowX setup, the tasks consist of: (1) put a carrot on the plate; (2) put an eggplant on the basket; (3) put a spoon on the towel; (4) stack a green cube on a yellow one.
We follow the standard evaluation protocol to test by randomizing both configurations of the environments. For the Google Robot tasks, we execute 300 trials of “Pick Coke Can”, 240 of “Move Near”, 216 of “Open/Close Drawer”, and 108 of “Place Apple in Closed Drawer”. For each WidowX task, we use 24 unique configurations. To ensure statistical significance, we test each configuration 10 times, yielding 240 rollouts per task. Reported results (Table 2) are the average success rates across these trials. Please refer to SIMPLER [33] for more details.
For a fair comparison, we adopt the published performance metrics for RT-1-X [13], Octo-base [54], OpenVLA [30], RoboVLMs [34], MoTo [11], and LAPA [67] directly from their respective papers. In the case of GR00T [53], we use the official pretrained checkpoint and performe fine-tuning on the RT-1/Bridge dataset following the authors’ published guidelines accordingly.
D.2 LIBERO
The LIBERO benchmark [40] evaluates knowledge transfer in multitask and lifelong robot learning problems for robotic manipulation, consisting of four task suites: LIBERO-Spatial, LIBERO-Goal, LIBERO-Object, LIBERO-Long. For fine-tuning of our models, we reinitialize the linear state encoder, action encoder, and action decoder, and tune the full parameters (except for the vision encoder). We fine-tune all the models for 50k gradient steps on the fine-tuning dataset on each task suite of LIBERO. We follow the standard evaluation protocol to test by randomizing the initial states of the environments. Reported results (Table 3) are the average success rates across on a total of 200 trials and a total of 20 trials per task on each task suite. Please refer to Section 4.2 and Appendix C of LIBERO [40] for more details.
Appendix E Real-world Robot Platforms Evaluation Details
E.1 Realman robot arm
The Realman robot arm setup is shown in Figure 8 (upper). We mount the gripper for Inspire Robot to the Realman RM75 robot arm. We use two camera views, including a primary view camera with the same view point as the images (used to demonstrate different tasks) shown in Figure 8 (upper) and a wrist camera. For fine-tuning of our models, we reinitialize the linear state encoder, action encoder, and action decoder, and tune the full parameters (except for the vision encoder). We fine-tune all the models for 60k gradient steps.
We collect data on the following five tasks with their task instructions:
  • Put-in: “Pick the green block from the table into the blue bowl”
  • Put-out: “Pick the green block from the blue bowl onto the table”
  • Push: “Push the green block to position X” where “X” indicates the nine positions written on the table.
  • Stack: “Stack the wooden block onto the green block”
  • Unstack: “Unstack the wooden block from the green block”
We collect 375 trajectories (75 trajectories for each task) for fine-tuning. The trajectories are collected at 10Hz. We post-process these trajectories to remove static frames with zero action, resulting in 120 steps on average in one trajectory.
We evaluate the fine-tuned model on seven groups with 10 trials for each group. The first five groups contain the tasks the same as data collection. The last two groups are designed to evaluate the generalization ability of the models. For the “change block color” group, we repeat the previous five tasks but change the green block into blue and red ones. For the “change table cover” group, we change the table cover from red to brown and blue ones.
The visualization example of each task for our model can be found in Figure 10.
Figure 10: Realman evaluation trajectory examples.
E.2 XHand dexterous hand
The Xhand setup is shown in Figure 8 (lower). The 12-dof Xhand is mounted on a 7-dof XArm robot arm. There are two camera views, including a main 3-rd view camera, and a wrist camera. During fine-tuning, we reinitialize linear encoder and decoder modules for both state and action to accommodate the hand’s higher dimensionality.
We use the dataset collected in [26] as our finetuning dataset, which comprises roughly 4,000 trajectories spanning 13 task categories and over 50 unique objects. For evaluation, we focus on five representative XHand tasks as depicted in Figure 8, namely pick-and-place, cube stacking, upright cup placement, water pouring, and ball flicking. Each task is assessed under “seen” and “unseen” conditions: in the seen setting, the same objects and backgrounds encountered during training are used, albeit with randomized tabletop positions and optional distractors; in the unseen setting, either the target objects or the scene background (or both) were never encountered during finetuning, totaling more than 20 novel objects. During evaluation, we conducted 50 evaluation runs for the pick-and-place task, 20 runs for cube stacking, and 10 runs for each of the remaining tasks. The visualization example of each task can be found in Figure 11 and Figure 12.
Figure 11: Xhand evaluation trajectory examples (part I).
Figure 12: Xhand evaluation trajectory examples (part II).

### v3

@@
 Xiaoyu Chen Hangxing Wei Pushi Zhang Chuheng Zhang Kaixin Wang Yanjiang Guo Tsinghua University Rushuai Yang Yucen Wang Xinquan Xiao Li Zhao Jianyu Chen Tsinghua University Jiang Bian Microsoft Research
-Visual-Language-Action (VLA) models have emerged as a popular paradigm for learning robot manipulation policies that can follow language instructions and generalize to novel scenarios.
-Recent work has begun to explore the incorporation of latent actions, an abstract representation of visual change between two frames, into VLA pre-training.
-In this paper, we introduce villa-X, a novel Visual-Language-Latent-Action (ViLLA) framework that advances latent action modeling for learning generalizable robot manipulation policies.
+Vision-Language-Action (VLA) models have emerged as a popular paradigm for learning robot manipulation policies that can follow language instructions and generalize to novel scenarios.
+Recent works have begun to explore the incorporation of latent actions, abstract representations of motion between two frames, into VLA pre-training.
+In this paper, we introduce villa-X, a novel Vision-Language-Latent-Action (ViLLA) framework that advances latent action modeling for learning generalizable robot manipulation policies.
 Our approach improves both how latent actions are learned and how they are incorporated into VLA pre-training.
-Together, these contributions enable villa-X to achieve superior performance across simulated environments including SIMPLER and LIBERO, as well as on two real-world robot setups including gripper and dexterous hand manipulation.
-We believe the ViLLA paradigm holds significant promise, and that our villa-X provides a strong foundation for future research.
-Keywords: Latent Action, Vision-Language-Action Model code: https://github.com/microsoft/villa-x site: https://aka.ms/villa-x
+We demonstrate that villa-X can generate latent action plans in a zero-shot fashion, even for unseen embodiments and open-vocabulary symbolic understanding.
+This capability enables villa-X to achieve superior performance across diverse simulation tasks in SIMPLER and on two real-world robotic setups involving both gripper and dexterous hand manipulation.
+These results establish villa-X as a principled and scalable paradigm for learning generalizable robot manipulation policies.
+We believe it provides a strong foundation for future research.
+Keywords: Latent Action, Vision-Language-Action Model
 1 Introduction
 Latent action learning has emerged as a promising approach for the pretraining of vision-language-action (VLA) models [12, 4, 31, 35, 10, 29, 70, 50], enabling learning from both robot data and human video data [68, 10, 50, 1].
-At the core of these methods is a Latent Action Model (LAM), which compresses visual changes between successive frames into latent tokens, referred to as latent actions.
-These latent actions are expected to capture motion semantics.
-These tokens function as pseudo-action labels for imitation learning and enriching robot policy training with abundant, action-free data.
-Yet two questions remain: how to learn latent actions and how to incorporate latent actions into VLA pre-training?
-In this paper, we introduce villa-X, a novel Visual-Language-Latent-Action (ViLLA) framework that advances both key aspects of latent action modeling.
-For the latent action learning component, a key limitation is the absence of direct alignment between visual transitions and underlying robot states and actions.
-Although robot data with low-level states and actions are often available alongside action-free videos during pretraining, current methods choose to ignore these signals and focus solely on the visual part.
-As a result, latent actions can remain ungrounded in the robot’s physical dynamics, resulting in a weak correspondence between latent action and robot behavior.
-To address this issue, we incorporate a proprio Forward Dynamics Model (FDM) module into the LAM as an auxiliary decoder.
-This module predicts future robot proprioceptive states and actions, encouraging the learned latent actions to better reflect the agent’s behavior—not just the observed visual changes.
-Moreover, this design improves the interpretability of the learned latent actions and enables them to be more easily translated into executable robot actions.
-As a result, latent actions become a robust intermediary between vision-based representations and low-level controls.
-To exploit this structure, we propose to model the latent action and robot action distribution jointly through a joint diffusion process, and conditioning robot action generation on latent actions generation through attention.
-Compared to existing methods on latent action pre-training, our approach allows for more effective and structured information transfer from latent actions to robot actions.
-We conduct comprehensive evaluations of villa-X on two simulated environments, SIMPLER and LIBERO, as well as two real-world setups, including various robot platforms with gripper and dexterous hand manipulation.
-Our results demonstrate that villa-X achieves superior performance, laying a robust foundation for future research in this domain.
-Our contributions are listed as follows:
-  • We improve latent action learning by introducing an extra proprio FDM, which aligns latent tokens with underlying robot states and actions and grounds them in physical dynamics.
+At the core of these methods is a Latent Action Model (LAM), which is designed to capture the motion semantics between successive frames into compact latent tokens.
+These tokens, referred to as latent actions, serve as pseudo-action labels, enriching robot policy training by enabling imitation learning on abundant, action-free data.
+While promising, the central challenge still lies in improving how latent actions can enhance robot policy learning.
+This motivates our investigation into two pivotal questions: how to better learn latent actions, and how to more effectively integrate them into VLA pre-training?
+In this paper, we introduce villa-X, a novel Vision-Language-Latent-Action (ViLLA) framework that advances both key aspects of latent action modeling.
+For the latent action learning component, existing latent action models [68, 10, 50, 1] typically compress latent actions based on visual signals, as shown in Figure 1 (a).
+However, while visual changes generally align with robot physical dynamics, certain motions, such as end-effector rotations or gripper movements, are subtle in pixel changes but critical for control.
+Vision-based models often pay less attention to these motions, a limitation also noted in recent work [9].
+As a result, the learned latent actions remain physically ungrounded, hindering effective knowledge transfer.
+To address this, we move beyond purely visual signals by leveraging structural cues for physical grounding.
+Specifically, we integrate a proprioceptive Forward Dynamics Model (proprio FDM) as an auxiliary decoder within our Latent Action Model (LAM), as shown in Figure 1 (b).
+This module predicts future robot proprioceptive states and actions by including embodiment context as inputs to help distinguish heterogeneous data.
+As a result, the latent actions become better grounded by focusing on visual changes that align with the agent’s physical dynamics.
+This makes the latent action a more effective bridge between vision and control, ultimately improving knowledge transfer.
+This framework is general and can be extended to other cues like end-effector keypoint detection or human hand pose estimation, which we leave for future work.
+To better leverage learned latent actions, we introduce a novel integration strategy for VLA pre-training.
+villa-X models latent and robot actions within a joint diffusion framework composed of two components: a latent action expert (ACT-latent) and a robot action expert (ACT-robot) as shown in Figure 2.
+Within this framework, an attention mechanism conditions robot action generation on latent actions generation.
+Compared to existing methods, this framework facilitates a more effective and structured transfer of information.
+We conduct comprehensive evaluations of villa-X across diverse environments.
+Our experiments yield two key findings.
+First, extensive ablation studies confirm that our proposed improvements to the latent action model and policy architecture outperform existing methods.
+Second, we show that by scaled pre-training, the latent action expert effectively plans for the future, and generalizes zero-shot to unseen embodiments and open-vocabulary symbolic icons.
+Collectively, villa-X achieves state-of-the-art performance in various tasks including numerous simulation tasks in SIMPLER and two real-world setups featuring various robotic platforms with both gripper and dexterous hand manipulators.
+This establishes a robust foundation for future research in the field.
+In summary, our main contributions are as follows:
+Figure 1: (a) A standard Latent Action Model (LAM) learns a latent action $z_{t}$ primarily through visual reconstruction, predicting a future frame $\hat{o}_{t+K}$ from the current frame $o_{t}$ and latent action $z_{t}$.
+(b) Our proposed model enhances this by adding a proprio-FDM.
+This auxiliary module predicts future robot states $\hat{q}_{t+1:t+K}$ and actions $\hat{a}_{t:t+K-1}$ conditioned on an embodiment context $c_{e}$, enabling the latent actions to be better grounded in physical dynamics.
+  • We improve latent action learning by introducing an extra proprio FDM, which grounds latent actions in physical dynamics by aligning with underlying robot states and actions.
   • We propose to jointly learn a latent action expert and a robot action expert through joint diffusion in the policy model, conditioning robot action prediction on latent actions to fully exploit their potential.
-  • Our method demonstrates superior performance on simulated environments as well as on real-world robotic tasks.
-The latent action expert can effectively plan into future with both visual and proprio state planning.
-Figure 1: (a) High-level architecture of our villa-X method. (b) Detailed structure of the Latent Action Model (LAM) component within villa-X.
+  • Through scaled pretraining, our latent action expert develops strong zero-shot generalization capabilities.
+This enables the effective transfer of knowledge across diverse simulated environments and real-world robotic tasks, leading to superior performance.
 2 Related Work
-Visual-Language-Action Models
+Vision-Language-Action Models
 Vision-Language-Action (VLA) models [4, 31, 50, 10, 29, 70, 35, 12] leverage pre-trained vision-language models (VLMs) to generate robot actions using visual and language cues.
@@
 Utilizing unlabeled trajectory data with pseudo-labels such as latent actions [68, 10], language sub-goals [47], or visual sub-goals [70] supports overcoming these challenges.
-Our method adopts the latent action approach and improves both the modeling of latent actions and their integration into VLA pretraining.
+Our method advances the latent action framework by enhancing both the modeling of latent actions and their integration into VLA pre-training.
 Modeling Latent Actions for VLA Pretraining
@@
 Dynamo [14] adopted a similar architecture, using inverse and forward dynamics models to shape state representations.
-LAOM [51] propose to use supervision to learn better latent actions in the presense of distractors on Mujoco environments.
 For robotic learning, methods have started to incorporate latent actions into VLA pretraining [68, 9, 10, 50, 1, 7, 66, 38].
 LAPA [68] proposes to learn from videos, and trains its latent actions and Vision-Language Model (VLM) using either human or robot video data.
-Concurrently, IGOR [10] learns latent actions from a mixture of human and robot videos, marking the first demonstration of successful latent action transfer between humans and robots in a unified action space for embodied AI.
+Concurrently, IGOR [9] learns latent actions from a mixture of human and robot videos, marking the first successful transfer of latent actions between humans and robots.
 Moto-GPT [10] co-fine-tunes both latent and robot action labels.
@@
 UniVLA [7] proposes a two-stage training pipeline to learn task-centric latent actions.
-More recent works like [66, 38] explore the continuous latent actions.
+More recent works like Yang et al.
+[66], Liang et al.
+[38] explore the continuous latent actions.
+[69] provides the analysis on the learned latent actions, while LAOM [49] uses supervision to learn latent actions that are robust to distractors in MuJoCo environments.
 By contrast, our approach jointly models latent and robot actions through a joint diffusion process, conditioning robot action generation on latent actions for more effective and structured information transfer.
-Our method offers better integration compared to LAPA [67] and GR00T [52], incorporates immediate visual context unlike Moto-GPT [11], and avoids inconsistencies caused by teacher forcing seen in Go-1 [1], achieving robust test-time reasoning.
+Our method improves upon prior work in several key aspects: it offers tighter integration than LAPA [68] and GR00T [50]; it incorporates immediate visual context, unlike Moto-GPT [10]; and it avoids teacher-forcing inconsistencies seen in Go-1 [1].
+These advantages collectively contribute to more robust reasoning at test time.
 3 Method
-Figure 2: Visualization of image pairs with similar latent actions.
-At a high level, our villa-X consists of two main components, as illustrated in Figure 1a:
+Figure 2: Architecture of ACT: A hierarchical policy that predicts latent action plans and conditions robot action generation on them, incorporating embodiment context and attention masking.
+Our method, villa-X, learns a physically grounded latent action space and uses it to train a VLA policy. The framework has two components:
   1. (i)
-    a LAM (Latent Action Model) module that infers latent actions from a pair of observations;
+    Latent Action Model (LAM) infers latent actions from a pair of observations, aligning these latent actions with robot dynamics via additional proprioceptive supervision.
   2. (ii)
-    a ACT (Actor) module that jointly models latent actions and robot actions given an initial visual observation and a textual task instruction.
-We describe each component in detail in the following subsections.
-Learning in villa-X is carried out in three stages: (1) pretraining the LAM module, (2) pretraining the ACT module, and (3) finetuning ACT on target-embodiment data.
-3.1 Latent Action Model
-Modeling latent actions enables the use of abundant action-free video demonstration data and also holds the promise of improving generalization between robot and human motions [67, 11].
-Building on these prior works, our LAM (depicted in Figure 1b) learns an inverse dynamics model (IDM) to extract latent actions $z_{t}$ from two adjacent video frames, $o_{t}$ and $o_{t+K}$.
-The interval $K$ is a configurable window size, chosen to balance perceptibility and generalizability: large enough to capture meaningful motion but small enough to maintain transferability.
-Previous works [67, 11] train IDM by jointly optimizing a forward dynamics model (FDM) to predict future frames.
-Given the current frame $o_{t}$ and a latent action $z_{t}$, the model minimizes the discrepancy between the predicted frame $\hat{o}_{t+K}$ and the ground truth $o_{t+K}$.
-Although robot data with low-level states and actions are often available alongside action-free video data, these approaches completely discards such information, focusing solely on visual changes.
-Since our ultimate objective is to use latent actions to facilitate real robots policy learning, bridging vision-language prompts and low-level controls would be beneficial.
-To this end, we propose augmenting LAM pretraining with proprioceptive supervision.
-Specifically, since $z_{t}$ represents a higher-level action abstraction, we introduce an additional proprio FDM module that predicts future robot states and actions $K$ steps ahead given the current robot state $q_{t}$ and the latent action $z_{t}$.
-This alignment of high-level abstractions and physical robot dynamics enhances the suitability of latent action for robot policy learning.
-However, because robot trajectories are often embodiment specific, conditioning solely on $q_{t}$ and $z_{t}$ may embed platform-dependent features in the latent action, hindering cross-embodiment transfer.
-To alleviate this issue, we provide the proprio FDM with an additional embodiment context input (e.g., dataset ID, control frequency), isolating transferable action representations from embodiment particulars.
-[Uncaptioned image] Implementation
-In practice, instead of predicting a single latent action between two frames (as Figure 1b depicts), LAM receives a sequence of $T_{\text{LAM}}$ observation frames as input and outputs $T_{\text{LAM}}-1$ latent actions between consecutive frames.
-The inverse dynamics model first encodes the frame sequence with a Spatial–Temporal Transformer [65], then applies value quantization to produce $T_{\text{LAM}}-1$ latent actions $z\in\mathbb{R}^{N_{z}\times D_{d}}$.
-The FDM is implemented using a Vision Transformer (ViT), while the proprio FDM is a two-layer MLP with dual output heads.
-The embodiment context input is embedded using learnable embeddings and concatenated with the robot state before being fed into the proprio FDM.
-For further implementation details and configuration specifics, please refer to the appendix.
-🖹 Training and Inference
-We train LAM on diverse data sources, including robot trajectories and human egocentric videos.
-For data lacking low-level state and action labels, we omit the proprio FDM branch and optimize only the FDM reconstruction loss.
-By leveraging this diverse data, our goal is to learn a generalizable intermediate representation for downstream policy learning.
-Once trained, only the IDM is required to extract latent actions from a pair of frames at inference, the FDM and proprioceptive FDM serve primarily as diagnostic visualization tools.
-We use a mixture of datasets containing both human video and robot action data including OpenX [13], Ego4D [22], etc.
-For more details, please refer to Appendix A.1.
-3.2 Actor Module
-Figure 3: Structure of the ACT module.
-In the same vein as common VLA models [30, 31, 5, 67, 52], our ACT module builds upon a pre-trained vision-language model (VLM) and is trained to model a generalizable robot control policy that outputs actions given a textual task instruction and an initial observation.
-With the pre-trained LAM, we are able to incorporate latent actions as an additional intermediate representation in our framework, rather than directly mapping vision and language inputs to low-level robot actions.
-The latent actions serve as a bridging language that connects task prompts to robot action planning.
-Figure 3 provides an illustration of our ACT module, which consists of three components: VLM, ACT-latent, and ACT-robot.
-Given a textual task prompt $l$ and a visual observation $o_{t}$, the VLM first extracts visual and language features.
-Then, the ACT-latent component, trained with a diffusion objective, models a sequence of $n$ latent actions $z_{t},z_{t+K},\ldots,z_{t+(n-1)K}$, conditioned on the features from the VLM.
-Subsequently, the ACT-robot component, also implemented as a diffusion model, predicts a sequence of $m$ low-level robot actions, conditioned on the same visual and language features as well as the latent actions.
-Conditioning across components is achieved using uni-directional attention over intermediate features, as indicated by the dashed line in Figure 3.
-Additionally, ACT-robot can incorporate wrist camera input when available.
-Unlike prior work [5], which feeds wrist camera data into the VLM from the start, our design is more computationally efficient.
-Another consideration is that we view the wrist camera input as more embodiment-specific and less generalizable, which is why we incorporate it into ACT-robot rather than the VLM.
-The first two components, VLM and ACT-latent, are designed to handle the more generalizable aspects of the policy.
-One might ask whether the length of the robot action sequence must match that of the latent action sequence, i.e., whether $m=nK$.
-In our design, this is not required.
-We allow different lengths for latent and robot action sequences, as long as they share the same starting time step $t$.
-For example, we may model 10 latent actions (corresponding to 20 robot actions if $K=2$) but choose to predict only 4 robot actions (effectively relying on the first 2 latent actions).
-This flexibility enables us to maximize the benefits of modeling latent actions.
-We highlight three key elements of our design for incorporating latent actions:
-  • Latent actions as a mid-level bridge.
-Previous works, including LAPA [67] and GR00T [52], treat latent actions similarly to robot actions and do not explicitly model the hierarchical structure between them.
-For example, GR00T simply treats latent actions as an additional action type and trains them alongside robot actions within a shared diffusion transformer.
-In contrast, our model treats latent actions as a distinct mid-level representation that bridges high-level vision and language prompts with low-level robot actions.
-  • Explicit transfer from latent to robot actions.
-Unlike LAPA [67] and GR00T [52], where the transfer from latent actions to robot actions occurs only implicitly through pre-trained weight initialization, our model enables an explicit connection.
-The robot action diffusion process is directly conditioned on the latent action diffusion process, allowing more effective and structured information transfer from latent actions to robot actions.
-  • Modeling latent actions sequences.
-While prior works (LAPA [67] and GO-1 [1] ) incorporate latent actions, they only model one-step predictions rather than planning over longer horizons at the latent action level.
-In contrast, our approach models a sequence of future latent actions, enabling structured planning at both the latent and robot action levels.
-Implementation  We use a pre-trained PaliGemma model [3] as our VLM.
-Both ACT-latent and ACT-robot adopt a transformer architecture following [69], and are trained using flow matching [39, 42].
-Wrist camera images are tokenized following the approach used in HPT [63].
-Similar to the proprio FDM in LAM, we incorporate embodiment context embeddings in ACT-robot to account for embodiment-specific information.
-Inspired by Moto [11] and RDT [43], we apply attention masking during training: with 50% probability, we completely disable attention from latent actions to robot actions; otherwise, we randomly mask out attention from 50% of the latent actions.
-Further implementation details are provided in the appendix.
-🖹 Training, Finetuning and Inference During training, we update all components of the model, including both the pre-trained and randomly initialized parts.
-The model is trained jointly using losses on both latent actions and robot actions, leveraging a mixture of datasets containing both human video and robot action data including OpenX [13], Ego4D [22], etc.
-Finetuning follows the same procedure as training, with the exception that, for unseen embodiments, we initialize a new context embedding.
-For robot action inference, we perform denoising steps simultaneously in both ACT-Latent and ACT-Robot.
-For more details, please refer to Appendix A.2.
+    ACTor Module (ACT) builds on a pre-trained vision-language backbone and jointly models sequences of latent and robot actions for improved planning and control.
+Training proceeds in three stages: (i) LAM pretraining on diverse datasets, (ii) ACT pretraining with joint latent–robot modeling, and (iii) embodiment-specific finetuning.
+3.1 Latent Action Model (LAM)
+Latent actions provide a compact intermediate representation, enabling the use of abundant human videos and improving cross-embodiment generalization [68, 10].
+Prior works typically learn a quantized codebook of latent actions via two modules: an Inverse Dynamics Model (IDM) and a visual Forward Dynamics Model (FDM).
+The IDM predicts a latent token $z_{t}$ from a frame pair $(o_{t},o_{t+K})$, while the FDM reconstructs the future observation $\hat{o}_{t+K}$ from $(o_{t},z_{t})$:
+$z_{t}=\text{IDM}(o_{t},o_{t+K}),\quad\hat{o}_{t+K}=\text{FDM}(o_{t},z_{t}).$ (1)
+This objective ensures consistency in visual change but ignores physical dynamics, producing latents that are insufficiently grounded when robot states are available.
+Proprioceptive Grounding.
+To address this, we introduce an additional proprioceptive Forward Dynamics Model (proprio-FDM) that predicts both future robot states and actions $K$ steps ahead, given the current state $q_{t}$ and latent $z_{t}$:
+$(\hat{q}_{t+1},...,\hat{q}_{t+K},\hat{a}_{t+1},...,\hat{a}_{t+K})=\text{proprio-FDM}(q_{t},z_{t},c_{e}),$ (2)
+where $c_{e}$ denotes an embodiment context described below. Optimizing visual and proprioceptive forecasting jointly encourages latent tokens to emphasize physical dynamics alongside visual changes.
+Disambiguating Heterogeneous Embodiments.
+Large-scale datasets mix embodiments with different morphologies and control frequencies.
+Naively conditioning the proprio-FDM on $(q_{t},z_{t})$ risks encoding embodiment-specific features into latents.
+We introduce a context vector $c_{e}$ comprising:
+$c_{e}=f(\text{dataset ID, control frequency}),$ (3)
+where dataset IDs are mapped to learnable embeddings, and frequencies are encoded using sinusoidal features passed through an MLP.
+These embeddings, concatenated with $q_{t}$, allow proprio-FDM to separate embodiment-specific dynamics while preserving latent action consistency across datasets.
+The full LAM thus optimizes image reconstruction loss, proprioceptive prediction loss, and vector-quantization commitments jointly.
+For human video lacking proprio labels, the proprio term is omitted.
+Finally, we adopt the continuous vector from the VQ codebook center as our latent actions.
+We refer readers to Appendix A for further implementation details.
+In summary, our LAM extends prior latent action models beyond compressing only visual changes to jointly modeling physical state transitions.
+While this work leverages robot proprioception for grounding, the framework is generic: alternative structural cues like end-effector keypoint detection or human hand pose estimation could replace low-level states, which we leave for future exploration.
+3.2 Actor Module (ACT)
+Our ACT module extends traditional vision-language-action (VLA) approaches by explicitly modeling both latent actions ($z_{t:t+(n-1)K}^{K}=(z_{t},z_{t+K}...,z_{t+(n-1)K}$)) and robot actions ($a_{t:t+m-1}=(a_{t},a_{t+1},...,a_{t+m-1})$) with a joint diffusion process.
+We factorize the policy into two conditional distributions:
+$\pi(a_{t:t+m-1},z_{t:t+(n-1)K}^{K}\mid o_{t},l,q_{t},c_{e})$ $=\underbrace{\pi_{\text{robot}}\bigl(a_{t:t+m-1}\mid z_{t:t+(n-1)K}^{K},\,o_{t},l,q_{t},c_{e}\bigr)}_{\text{ACT-robot}}$ (4)
+$\quad\cdot\;\underbrace{\pi_{\text{latent}}\bigl(z_{t:t+(n-1)K}^{K}\mid o_{t},\,l\bigr)}_{\text{ACT-latent}}.$
+where $o_{t}$ is the observation, $l$ the language instruction, $q_{t}$ the proprioceptive state, and $c_{e}$ the embodiment context.
+Additionally, the low-level policy can optionally incorporate wrist camera input if available.
+This explicit modeling and factorization improves upon prior methods, such as LAPA [68], which rely on latent actions only through pre-trained weight initialization.
+By contrast, our method treats latent actions as a distinct mid-level representation that bridges high-level vision and language prompts with low-level robot actions, and allow for allowing more effective and structured information transfer from latent actions to robot actions.
+Architecture.
+ACT (Figure  2) comprises three experts with a blockwise causal attention mask:
+  • VLM: Encodes the visual-language inputs into high-level features.
+  • ACT-latent: Latent action expert that predicts latent action tokens for mid-level planning, conditioning on VLM features.
+  • ACT-robot: Robot action expert that produces the low-level action chunk, conditioning on VLM features, predicted latents and additional control-specific inputs including proprioceptive states and embodiment context.
+Attention Masking Strategies.
+A key aspect of ACT is how we maintain a robust dependence on the latent tokens without letting the policy learn trivial shortcuts.
+Inspired by Moto [10] and RDT [41], we adopted the stochastic Masking strategy.
+During training, we stochastically mask the attention from robot actions to latent actions.
+In 50% of the cases, all robot-to-latent attention are masked; otherwise, 50% of the latent tokens are randomly masked.
+This can prevent the robot-action branch from overly relying on latent actions and leading to improved robustness.
+We found this design crucial in practice.
+Joint Diffusion Modeling.
+For implementation, villa-X models the joint distribution of the future robot actions $a_{t:t+m-1}$ and latent actions $z_{t:t+(n-1)K}^{K}$ using a conditional flow matching framework.
+For notational simplicity, we group these actions into a single variable $x_{t}$ and denote the conditioning inputs $(o_{t},l,q_{t},c_{e})$ as $O_{t}$.
+The objective is to train a network $v_{\tau}^{\theta}$ that minimizes the flow matching loss:
+$L_{\tau}(\theta)=\mathbb{E}_{p(x_{t}\mid O_{t}),\,q(x_{t}^{\tau}\mid x_{t})}\left\|v_{\tau}^{\theta}(x_{t}^{\tau},O_{t})-u(x_{t}^{\tau}\mid x_{t})\right\|^{2}$ (5)
+where $\tau\in[0,1]$ denotes flow matching timestep.
+In practice, we first sample random noise $\epsilon\sim N(0,I)$ to create a noisy target $x^{\tau}_{t}=\tau x_{t}+(1-\tau)\epsilon$.
+The network $v_{\tau}^{\theta}(x_{t}^{\tau},O_{t})$ is then trained to predicted the denoising vector field $u(x_{t}^{\tau}\mid x_{t})=\epsilon-x_{t}$.
+During training, we sample $\tau$ from beta distributions.
+Notably, the explicit factorization in Eq.
+4 is achieved by block-wise causal attention.
 4 Experiments
@@
   • Does our improved LAM learn higher-quality latent actions?
-  • Can ACT-latent successfully plan future motions?
   • Can the actor module effectively leverage the pre-trained latent actions?
+  • By scaling pre-training, can the latent actor module effectively plans for the future and generalize zero-shot to unseen embodiments and open-vocabulary concepts in symbolic icons?
   • How does villa-X compare to existing VLA baselines in both simulated benchmarks and real-world robot tasks?
@@
 To assess its impact, we compare our model (denoted w/pp) to a variant without the proprio FDM module (denoted wo/pp).
+Figure 3: Probing experiment results.
 Probing  First, a core expectation for latent actions is that they should carry information useful for predicting low-level robot actions.
@@
 We train the MLP on the training split of LIBERO and evaluate it using the L1 loss on the validation split.
-Figure 4: Probing experiment results.
 We are interested in how closely the predicted actions match the ground truth.
-In LIBERO, the robot action space has 8 dimensions: 3 for position, 4 for rotation, and 1 for the gripper.
+In LIBERO, the robot action space has eight dimensions: three for position, four for rotation, and one for the gripper.
 Rather than averaging the error across dimensions, we focus on the maximum L1 error across all action dimensions, as we want to avoid large deviations in any single aspect of the action.
@@
 This demonstrates the effectiveness of the proprio FDM module in capturing information from the robot actions.
+We further visualize the learned latent actions, and perform more ablations on LAM.
+Please refer to Appendix D for more details.
 Table 1: Evaluation results on SIMPLER for different variants of our villa-X(top group) and alternative approaches for incorporating latent actions (bottom group).
-”Ours” refers to the w/pp described in the main text.
-Method Google robot WidowX robot Pick Move Drawer Place Avg.
+“Ours” refers to the w/pp described in the main text.
+Method Google robot WidowX robot Pick Move Drawer Avg.
 Carrot Eggplant Spoon Cube Avg.
-Ours 81.7 55.4 38.4 4.2 44.9 24.2 71.7 48.3 19.2 40.8 wo/pp 77.0 52.7 42.6 2.8 43.8 22.5 57.5 43.3 5.9 32.3 wo/LAM 42.1 24.6 38.4 0.0 26.3 25.8 60.8 36.7 9.2 33.1 LAPA-style 64.7 28.8 38.0 5.6 34.2 0.8 0.0 2.5 0.8 1.0 Go-1-style 29.0 38.0 31.3 4.6 25.7 5.8 50.8 1.7 1.0 14.8
+Ours 81.7 55.4 38.4 58.5 24.2 71.7 48.3 19.2 40.8 wo/pp 77.0 52.7 42.6 57.4 22.5 57.5 43.3 5.9 32.3 wo/LAM 42.1 24.6 38.4 35.0 25.8 60.8 36.7 9.2 33.1 LAPA-style 64.7 28.8 38.0 43.8 0.8 0.0 2.5 0.8 1.0 Go-1-style 29.0 38.0 31.3 32.8 5.8 50.8 1.7 1.0 14.8
 Policy Pre-training  Next, we compare how the latent actions generated by different variants of LAM (w/pp and wo/pp) influence policy pre-training.
@@
 The resulting policies are evaluated in the SIMPLER environment [32], a simulation benchmark explicitly designed to mitigate the gap between simulated and real‐world robotic environments.
-It comprises two platforms, namely Google robot and WidowX robot, each with four manipulation tasks.
-We evaluate our method on all the 8 tasks on the visual matching setting.
+It comprises two platforms: the Google robot with three manipulation tasks and the WidowX robot with four.
+We evaluate our method on the visual matching setting.
 The results are summarized in Table 1.
@@
 The performance of wo/LAM is significantly worse, indicating that pre-training with latent actions is essential.
-LAM visualization
-Figure 2 visualizes image pairs sharing the same latent action, demonstrating that these pairs correspond to similar underlying robot behaviors.
-To further demonstrate the transfer ability of our LAM, we extract latent actions from arbitrary video sequences, map them to robot actions using the proprio FDM, and execute the resulting robot action in SIMPLER simulator.
-The simulated motions closely reproduce the original demonstrations, indicating that latent actions are both aligned with and grounded in the robot’s actions.
-For more examples, please check Appendix C.1.
-Figure 5: Transfer of human demonstrations to new scenes via latent actions and a world model. (a) moving right. (b) moving downwards. Please visit our website for more demos.
-Figure 6: Transfer of video demonstrations into robot action executions via latent actions and a proprio FDM. Please visit our website for more demos.
-LAM transfer consistency
-We provide two visualizations to further evaluate the transfer consistency of the learned latent actions.
-Specifically, we first use the trained LAM to extract a sequence of latent actions from a source video.
-These latent actions are then applied to different initial images, allowing a world model to generate subsequent frames conditioned on these actions.
-As shown in Figure 5, the transfer rollouts (highlighted in blue) successfully identify the movable parts and demonstrate behaviors consistent with the original video.
-We also evaluate the alignment between latent and robot actions.
-To do this, we use the proprio FDM to decode latent actions into executable robot actions.
-These actions are then executed in the SIMPLER simulator, with the resulting rollouts visualized in Figure 6.
-In each example, the top row shows the original video demonstration from which latent actions were extracted, while the bottom row displays the corresponding simulated execution.
-The simulated motions closely match the original demonstrations, confirming that the latent actions learned by villa-X are well-aligned with and grounded in the robot actions.
-4.2 Can ACT-latent successfully plan future motions?
-Figure 7: Transfer of video demonstrations into robot action executions via latent actions and a proprio FDM. Please visit our website for more demos.
-In this experiment, we demostrate the motion planning capabilities of ACT-latent by visualizing its planned actions with a world model.
-For a given initial image and language instruction, ACT-latent generates a sequence of latent actions, and the world model takes the inital frame and latent actions as inputs to render the latent actions into planned future videos.
-Figure 7 shows results for both in-distribution and out-of-distribution samples, where the in-distribution samples are randomly sampled from the validation set but from the same datasets as used in training, and the out-of-distribution samples are from new real-world scenarios.
-The results demonstrate that ACT-latent successfully follows the language instructions to solve the tasks, while accurately identifying target objects and generating latent actions that successfully follow the instructions.
-ACT-latent also successfully identifies the concepts in the emoji, which rarely appears in robot datasets, suggesting that villa-X keeps the general vision-language capabilities in the initial VLM model after pre-training.
 4.2 Can the actor module effectively leverage the pre-trained latent actions?
@@
 Compared to other two approaches, our method achieves significantly higher performance, validating the effectiveness of our design for incorporating latent actions into VLA pre-training.
+More ablation studies on policy designs can be found in Appendix F.
+4.3 Zero-shot Generalization for Latent Actor Module
+Figure 4: Visualization of zero-shot latent plans on an unseen embodiment. Each pair of images shows the starting frame (left) and the ending frame (right), with the instruction displayed above.
+To evaluate the zero-shot generalization capabilities of ACT-latent in planning, we conducted a real-world visualization experiment focused on its ability to handle new embodiments and understand novel open-vocabulary symbols.
+For embodiment generalization, we used a Realman robot arm, a new embodiment never seen during training.
+To assess open-vocabulary generalization, we designed a set of symbol cards, testing the model’s ability to comprehend concepts typically absent from standard robotics datasets.
+The evaluation process is as follows: given a starting image and a language instruction (e.g., “touch the corn”), ACT-latent first generates a sequence of latent actions.
+Then a separately trained world model is used to render this sequence into a video, allowing us to verify the effectiveness of the plan.
+As shown in Figure 4, the rendered trajectories confirm that the model successfully generates latent plans that follow the commands. These results highlight two key capabilities of our approach:
+  • Embodiment Generalization: ACT-latent successfully identifies and controls this unseen robot arm, indicating that its learned knowledge is embodiment-agnostic and readily transferable to the new robotic platform.
+  • Open-Vocabulary Understanding: The model’s ability to interact correctly with symbol concepts reveals that villa-X retains the general-purpose vision-language capabilities of the original Vision-Language Model (VLM) after pre-training.
+More visualizations can be found in Appendix E. To evaluate how effectively villa-X uses this knowledge, we measure its success rates on a variety of control tasks in the following sections.
 4.4 Evaluating villa-X in Simulation
-4.4.1 SIMPLER Benchmark
 Table 2: Comparison on SIMPLER of villa-X and existing methods.
 Methods marked with * are evaluated directly after pretraining, whereas other methods are evaluated after post-training on corresponding dataset.
-Method Google Robot WidowX Robot Pick Move Drawer Place Avg.
+Method Google Robot WidowX Robot Pick Move Drawer Avg.
 Carrot Eggplant Spoon Cube Avg..
-RT-1-X * 56.7 31.7 59.7 21.3 42.4 4.2 0.0 0.0 0.0 1.1 Octo-base * 17.0 4.2 22.7 0.0 11.0 8.3 43.1 12.5 0.0 16.0 OpenVLA * 16.3 46.2 35.6 0.0 24.5 0.0 4.1 0.0 0.0 1.0 RoboVLMs * 72.7 66.3 26.8 36.1 50.5 25.0 0.0 20.8 8.3 13.5 RoboVLMs 77.3 61.7 43.5 24.1 51.7 20.8 79.2 45.8 4.2 37.5 GR00T-N1.5 69.3 68.7 35.8 4.0 44.5 54.3 61.3 75.3 57.0 62.0 MoTo 74.0 60.4 43.1 N/A N/A N/A N/A N/A N/A N/A LAPA N/A N/A N/A N/A N/A 45.8 58.3 70.8 54.2 57.3 Ours w/o latent 56.3 25.8 27.3 13.9 30.8 31.3 74.6 61.7 28.3 49.0 Ours 98.7 75.0 59.3 5.6 59.6 46.3 64.6 77.9 61.3 62.5
+RT-1-X * 56.7 31.7 59.7 49.4 4.2 0.0 0.0 0.0 1.1 Octo-base * 17.0 4.2 22.7 14.6 8.3 43.1 12.5 0.0 16.0 OpenVLA * 16.3 46.2 35.6 32.7 0.0 4.1 0.0 0.0 1.0 RoboVLMs * 72.7 66.3 26.8 55.3 25.0 0.0 20.8 8.3 13.5 RoboVLMs 77.3 61.7 43.5 60.8 20.8 79.2 45.8 4.2 37.5 $\pi_{0}$ 72.7 65.3 38.3 58.7 0.0 62.5 29.1 16.6 27.1 $\pi_{0}$-FAST 75.3 67.5 42.9 61.9 21.9 66.6 29.1 10.8 32.1 OpenVLA-OFT 72.3 69.6 47.2 63.0 4.2 N/A 12.5 8.3 N/A GR00T-N1.5 69.3 68.7 35.8 57.9 54.3 61.3 75.3 57.0 62.0 TraceVLA 45.0 63.8 63.1 57.3 16.6 65.0 12.5 16.6 27.7 Magma 75.0 53.0 58.9 62.3 29.2 91.7 37.5 20.8 44.8 MoTo 74.0 60.4 43.1 59.2 N/A N/A N/A N/A N/A LAPA N/A N/A N/A N/A 45.8 58.3 70.8 54.2 57.3 Ours w/o latent 56.3 25.8 27.3 36.5 31.3 74.6 61.7 28.3 49.0 Ours 98.7 75.0 59.3 77.7 46.3 64.6 77.9 61.3 62.5
 Baselines and Experimental Setup
-We use the SIMPLER benchmark as described above. In this section, we compare against two categories of prior work:
-  • Vision-Language-Action (VLA) models: RT-1-X [13], Octo-base [54], OpenVLA [30], RoboVLMs [34], which learn policies solely from mixed robot datasets.
+We use the SIMPLER benchmark as described above. In this section, we compare against several categories of prior work:
+  • Vision-Language-Action (VLA) models: RT-1-X [12], Octo-base [52], OpenVLA [29], RoboVLMs [34], $\pi_{0}$ [4], $\pi_{0}$-FAST [54], OpenVLA-OFT [30], which learn policies solely from mixed robot datasets.
   • Joint policy learning and world modeling method: GR00T-N1.5  [51], which aligns the model with target future embeddings.
+  • Visual trace methods: TraceVLA [71], Magma [67], which learn planning on the extract visual traces from videos.
   • Latent-Action based methods: MoTo [10] and LAPA [68], which additionally exploit unlabelled videos by inferring latent actions.
 Except where noted (*), all models follow a two‐stage pretraining–finetuning protocol, including a general pretraining phase on large‐scale mixed data, followed by finetuning on a dataset of specific embodiment.
-We also include an ablation (villa-X w/o latent) that removes our latent‐action expert while keeping all other components unchanged.
-Reported baseline scores are taken directly from the original publications, while missing entries are marked as N/A.
+We also include an ablation (Ours w/o latent) that removes our latent‐action expert while keeping all other components unchanged.
+Baseline scores are cited from their original publications or other relevant literature, while missing entries are marked as N/A.
 Experimental Results
 Table 2 summarizes the success rates on both platforms.
-Our full model achieves the highest score on average success rate on the Google robot (59.659.6%) and the WidowX robot (62.562.5%).
-This improvement over VLA methods, which cannot exploit unlabelled video, demonstrates the benefit of our incorporating human videos into policy learning.
-Moreover, our approach outperforms other latent‐action methods, indicating that our specific mechanism for leveraging video data is more effective.
+Our full model achieves the highest score on average success rate on the Google robot (77.777.7%) and the WidowX robot (62.562.5%).
+This improvement over VLA methods, which cannot exploit unlabelled video, demonstrates the benefit of incorporating human videos into policy learning.
+Moreover, our approach outperforms other video learning and latent‐action methods, indicating that our specific mechanism for leveraging video data is more effective.
 Finally, the gap between our full model and the “villa-X w/o latent” ablation confirms that the latent‐action expert is essential for achieving these gains.
-4.4.2 LIBERO Benchmark
-The LIBERO benchmark [40] evaluates knowledge transfer in multitask and lifelong robot learning problems for robotic manipulation, consisting of four task suites: LIBERO-Spatial evaluates the model’s performance under novel layouts with the same task and object types, LIBERO-Goal evaluates the model’s performance under novel tasks with the same object types and layouts, LIBERO-Object evaluates the model’s performance under novel object types with the same tasks and layouts, LIBERO-Long evaluates the model’s performance under diverse set of objects, layouts and backgrounds.
-Each task suite contains 10 tasks with 50 human demonstrations per task for fine-tuning.
-Baselines and Experimental Setup
-We compare with the following existing models: Diffusion Policy [12] trained from scratch, Octo [54], and OpenVLA [30].
-All models follow a two-stage pretraining-finetuning protocol.
-We finetune villa-X and villa-X w/o latent on the demonstration data of the each task suite separately, and test on the LIBERO simulator for 10 tasks and 20 trials per task on each task suite.
-Experimental Results
-Table 3 summarizes the success rates on each task suite of LIBERO.
-Our model achieves better performance than existing methods in all the four task suites.
-Also, our model with latent action achieves higher performance on 3 of the 4 task suites and average performance, confirming that the proposed latent action expert improves the manipulation performance.
-Table 3: Evaluation on 4 LIBERO task suites of villa-X and existing methods.
-Method Spatial Object Goal Long Average Diffusion Policy [12] 78.3 92.5 68.3 50.5 72.4 Octo-base [54] 78.9 85.7 84.6 51.1 75.1 OpenVLA [30] 84.7 88.4 79.2 53.7 76.5 Ours w/o latent 86.0 86.5 85.0 70.0 81.9 Ours 97.5 97.0 91.5 74.5 90.1
 4.5 Evaluating villa-X on Real-world Robots
+To assess real-world generalization, we deploy villa-X on two platforms: a Realman arm with a gripper and an XArm with a 12-DoF XHand as shown in Figure 5.
+Realman robot arm with gripper
+We use a 6-DoF Realman RM 75 with a 1-DoF Inspire gripper, fine-tuning and evaluating on five tasks: Pick-in (pick the block into a bowl), Pick-out (pick the block out of a bowl), Stack (stack the block onto another block), Unstack (unstack the block from another block), and Push (push the block to a given location).
+The fine-tuning set contains 375 teleoperated trajectories (75 per task); the object layout and table are fixed, while object positions vary.
+We conduct two sets of evaluation: In task evaluation, we remain the table setup the same as data collection; in generalization evaluation, we change the color of the block and table cover.
+For each task, we conduct 10 trials with distinct object positions; positions and lighting are identical across policies.
+As shown in Table 4, villa-X outperforms all baselines in both settings.
 Figure 5: Real-world robot evaluation platforms: (top) Realman robot arm platform with a gripper and (bottom) Xarm robot arm with Xhand dexterous hand.
 Platform setups are shown on the left, with corresponding evaluation tasks on the right.
-Please visit our website for more videos.
 Table 3: Evaluation on Xarm robot arm of villa-X and existing methods.
@@
 Method Pick in Pick out Push Stack Unstack Change block color Change table cover GR00T 30 70 10 10 60 50 30 Ours w/o latent 40 80 30 60 70 40 30 Ours 30 100 50 50 100 60 60
-To further assess generalization, we evaluate our approach on two real‑world robotic platforms: a Realman robot arm with a gripper, and an Xarm manipulator equipped with a 12‑DOF Xhand dexterous hand.
-See Figure 8 for details.
-Realman robot arm with gripper
-On the robot-arm manipulation platform, we use a 6-DoF Realman RM 75 robot arm and a 1-DoF Inspire gripper.
-We fine-tune and evaluate our policy on “Pick-in” (pick the block into a bowl), “Pick-out” (pick the block out of a bowl), “Stack” (stack the block onto another block), “Unstack” (unstack the block from another block), and “Push” (push the block to a given location) tasks.
-For fine-tuning, we collect a dataset of 375 trajectories through teleoperation (with 75 trajectories for each task), where the object layout and table setup are fixed and only the object locations are dynamic.
-We conduct two sets of evaluation: In task evaluation, we remain the table setup the same as data collection; in generalization evaluation, we change the color of the block and table cover.
-We evaluate the policy with 10 trials per task with each trial conducted with different object location, where all the experiment settings are the same for the evaluation of different policies (including the object location and lightening condition).
-The results are shown in Table 5, demonstrating that villa-X outperforms existing baselines in both task and generalization evaluation sets.
 Xarm robot arm with Xhand dexterous hand
@@
 In this paper, we presented villa-X, a novel Visual-Language-Latent-Action (ViLLA) framework that improves both the learning of latent actions and their incorporation into VLA pre-training.
-Our experiments demonstrate that our enhanced Latent Action Model learns higher-quality latent actions, and our improved policy model more effectively leverages these learned actions.
+Our experiments demonstrate that our enhanced Latent Action Model learns higher-quality latent actions, and our improved policy model more effectively leverages these learned latents.
+The learned latent action expert can even generalize zero-shot to an unseen embodiment, showing strong generalization ability.
 Overall, our method exhibits superior performance in both simulated environments and real-world robotic tasks.
@@
 We leave this aspect as future work to further improve the capability of the ViLLA framework.
-Appendix A Model Design
-A.1 Latent Action Model
-There are four modules in our latent action model.
-A ST-Transformer based inverse dynamic model (IDM), a vector quantization module, an image reconstruction forward dynamic model (FDM) and a robot proprio FDM.
-The IDM takes an 8 × 224 × 224 video clip as input, following a Patch Embedding with a patch size of 14 and 12 ST-blocks, where each block has a hidden dimension of 768 and 32 attention heads.
-The codebook size of vector quantization is set to 32.
-Image reconstruction FDM is a 12-layer ViT-base network, and robot proprio FDM is a 2-layer MLP with a hidden size of 768.
-Our model is trained with a learning rate of 1.5e-4, a batch size of 512, and a 2000-step linear warmup.
-We use the same loss weight for the image FDM and proprio FDM.
-The pretraining takes 4 days on 128 NVIDIA A100 GPUs.
-A.2 Actor Module
+Appendix A Additional Implementation Details for LAM
+In this appendix, we provide extended information on the architecture, training protocols, and inference behavior for our Latent Action Model (LAM).
+A.1 Architecture Overview
+Our LAM comprises four main modules:
+  1. (i)
+    Spatial–Temporal Transformer (ST-Transformer) Inverse Dynamics Model (IDM): Takes a video clip (by default, $8\times 224\times 224$) as input.
+We employ patch embedding with a patch size of $14$ and stack $12$ ST-blocks [65], each with a hidden dimension of $768$ and $32$ attention heads.
+  2. (ii)
+    Vector Quantization (VQ) Module: Maps the continuous IDM outputs to discrete latent tokens, each associated with a codebook entry.
+We set the codebook size to $32$.
+While the model internally uses discrete token indices during training, the continuous codebook centers are used in downstream modules.
+  3. (iii)
+    Image Reconstruction Forward Dynamics Model (FDM): A $12$-layer Vision Transformer (ViT)-base network that takes the current frame $o_{t}$ and a latent action $z_{t}$ to predict $\hat{o}_{t+K}$.
+  4. (iv)
+    Proprioceptive Forward Dynamics Model (proprio-FDM): A 2-layer MLP with dual output heads to predict future robot states $\hat{q}_{t+i}$ and low-level robot actions $\hat{a}_{t+i}$.
+This module takes the current robot state $q_{t}$, the latent $z_{t}$, and an embodiment context vector $\mathbf{c}_{e}$.
+Rather than predicting a single latent action per pair of frames, the ST-Transformer-based IDM processes a sequence of $T_{\text{LAM}}$ frames, resulting in $T_{\text{LAM}}-1$ latent tokens.
+We use $T_{\text{LAM}}=8$.
+By reconstructing future frames with the FDM and future states/actions with the proprio-FDM, the model learns a latent representation that is both visually and physically grounded.
+A.2 Training Details
+We train our LAM on a combination of human egocentric data (e.g., Ego4D [21]) and robot trajectories (e.g., OpenX [12]).
+Samples lacking low-level robot annotations (e.g., human videos) exclude the proprio-FDM branch, using only the visual FDM objective.
+Hyperparameters.
+We use a batch size of $512$ and a learning rate of $1.5\times 10^{-4}$, with a $2000$-step linear warmup.
+Training lasts approximately $4$ days on $128$ NVIDIA A100 GPUs.
+Both the visual FDM and proprio-FDM share the same weighting in the overall loss.
+Throughout training, each latent token is discretized via the VQ module but is represented by its continuous codebook center in subsequent network components.
+A.3 Inference Behavior and Diagnostics
+During inference, only the IDM is required to extract latent tokens from consecutive frames.
+The FDM and proprio-FDM are typically retained for diagnostic and visualization purposes, allowing us to examine whether the learned latent tokens accurately capture future frame content, robot states, and actions.
+This reconstruction-based analysis aids in understanding and debugging the physical grounding of the latent representation.
+Appendix B Additional Implementation Details for Actor Module
 Our VLA model comprises three components.
@@
 This combined masking strategy encourages the model to learn robust, generalizable policy that will balance the predicted latent actions as well as the input image and instruction.
+During training, we sample $\tau$ from different beta distributions for latent actions and robot actions, which biases the timesteps for latent actions towards the noisier regime.
 Each expert contains approximately 300 M parameters and is trained from scratch.
@@
 The pretraining takes 4 days on 64 NVIDIA A100 GPUs.
-Appendix B Dataset
-B.1 Data Mixture
+Appendix C Dataset
+C.1 Data Mixture
 We curated a data mixture by combining both robot data and action-free human videos for our pretraining phase.
@@
 For OpenX dataset, our base data mixture is created primarily based on [29, 52].
-In total, we use 223.5M trajectories with 1.6M frames of robot data.
+In total, we use 1.6M trajectories with 223.5M frames of robot data.
 For human videos, we use a mixture of Ego4D [21], EgoPAT3D [37], EGTEA Gaze+ [36], EPIC-KITCHENS [15], HO-Cap [62], HOI4D [42], HoloAssist [64], RH20T [18], Something Something V2 [20].
@@
 A full breakdown of our data mixture is listed in Table 5.
-B.2 Data Preprocessing
+C.2 Data Preprocessing
 For data cleaning, we adopt EgoHOD [53], a curated subset of Ego4D [21], and further filter the videos based on visual quality to ensure high-quality inputs for training.
@@
 Table 5: Our training data mixture used during the pretraining phase.
-Appendix C LAM visualization results
-C.1 Image Pairs with Similar Latent Actions
+Appendix D LAM visualization and More Ablations
+D.1 Image Pairs with Similar Latent Actions
 Figure 6: Visualization of image pairs with similar latent actions.
-Figure 9 visualizes additional image pairs sharing the same latent action, demonstrating that these pairs correspond to similar underlying robot behaviors.
+Figure 6 visualizes image pairs sharing the same latent action, demonstrating that these pairs correspond to similar underlying robot behaviors.
 The results demonstrate that similar latent actions represent the similar robot behaviors and low-level actions, in regardless of which embodiment (including human and different robots) is executing such action.
 This results support that villa-X learns cross-embodiment prior knowledge for manipulations with latent actions.
-Appendix D Simulation Evaluation Details
-D.1 SIMPLER Benchmark
+D.2 Transfer Video Demonstrations into Robot Actions through LAM and Proprio FDM
+Figure 7: Transfer robot video demonstrations into robot actions through LAM and proprio FDM in SIMPLER simulator.
+Upper: the SIMPLER rollout closely reproduce the motion of moving downwards, Bottom: the SIMPLER rollout closely reproduce the motion of moving right.
+Figure 8: Transfer human video demonstrations into robot actions through LAM and proprio FDM in SIMPLER simulator.
+Upper: the SIMPLER rollout closely reproduce the motion of moving right; Middle: the SIMPLER rollout closely reproduce the motion of moving forward and backward; Bottom: the SIMPLER rollout closely reproduce the motion of moving right.
+To further demonstrate the transfer ability of our LAM, we extract latent actions from videos of task demonstrations, map them to robot actions using the proprio FDM, and execute the resulting robot actions in the SIMPLER simulator.
+The results are presented in Figure 7 and Figure 8.
+In each figure, the top row shows the video demonstrations used by LAM to extract latent actions, while the bottom row displays the corresponding SIMPLER simulation results, where real actions decoded from the latent actions using proprioceptive FDM are executed.
+Specifically, Figure 7 illustrates robot-to-robot transfer, and Figure 8 illustrates human-to-robot transfer.
+The simulated motions closely reproduce the original demonstrations, indicating that latent actions learned by villa-X are both aligned with and grounded in the robot’s actions.
+D.3 More Ablations on LAM
+To validate the contribution of the embodiment context in our proprio-FDM, we further conducted an ablation study comparing our full method (”Ours”) against a version without the context (”Ours w/o context”).
+Both models were trained on 10
+(1) Performance on validation dataset: We measured the reconstruction loss of visual FDM and proprio FDM on the validation set:
+Method Visual FDM loss ($\downarrow$) Proprio FDM loss ($\downarrow$) Ours w/o context 0.068 0.078 Ours 0.057 0.070 Relative improvement 16.2% 10.3%
+Table 6: Performance comparison on the validation set.
+(2) Zero-Shot Generalization to a Novel Embodiment: We evaluated the model on our dataset collected on our Realman robot arm dataset (from Section 4.4), an embodiment completely unseen during training.
+We then conducted the action probing experiment described in Section 4.1 by inferring latent actions with IDM and training a new MLP to predict robot actions from the latent actions.
+Method Probing loss ($\downarrow$) Probing loss (xyz) ($\downarrow$) Probing loss (rot) ($\downarrow$) Probing loss (gripper) ($\downarrow$) Ours w/o context 0.165 0.0675 0.00861 0.928 Ours 0.152 0.0574 0.00619 0.873 Relative improvement 7.9% 15.0% 28.1% 5.9%
+Table 7: Zero-shot generalization to an unseen embodiment.
+The results from both experiments demonstrate that the embodiment context improves performance and aids generalization to novel embodiments.
+We hypothesize that while the visual FDM provides general transferability by aligning latent actions with visual changes, the proprio-FDM grounds these latent actions in robot physical dynamics.
+However, due to data heterogeneity (e.g., different action definitions / controllers, as discussed previously), the model requires the embodiment context to disambiguate different embodiments and learn a more consistent, grounded latent action space.
+Appendix E Latent Action Expert Visualization
+Figure 9: Generated image sequence jointly by the latent expert and the world model via latent actions, following different instructions from the same initial image (Part I).
+Figure 10: Generated image sequence jointly by the latent expert and the world model via latent actions, following different instructions from the same initial image (Part II).
+In this experiment, we demonstrate the performance of the latent expert by passing its prediction through the image reconstruction FDM that takes the latent action as inputs and predicts the future observations, which forms a simulated environment for the iteratively executing the latent expert.
+Starting from a single initial image, the latent expert and image reconstruction FDM jointly generate different behaviors in videos that follow diverse instructions using only latent actions.
+We experiment with initial images from RT-1 and Bridge dataset, and show the image clips of generated videos in Figure 9 and Figure 10 with different language instructions.
+The results show that the latent expert properly follows the language instructions for task solving, where the latent expert properly recognizes the target objects and predict latent actions that move towards the target object.
+Appendix F More Ablations on policy model
+We primarily conducted ablation studies on two main components: (1) the attention mask and (2) the embodiment context.
+Our experiments follow the same setting as Table 1 in the main paper.
+The ablation results below show that both the attention mask and embodiment context are effective in improving performance on two robot platforms: Google Robot and WidowX Robot.
+Table 8: Ablation study results for the policy model.
+The first columns (Pick, Move, Drawer, Place) refer to the Google Robot, and the last columns (Carrot, Eggplant, Spoon, Cube) refer to the WidowX Robot.
+All numbers are success rates (%).
+Method Pick Move Drawer Avg.
+Carrot Eggplant Spoon Cube Avg.
+Ours 81.7 55.4 38.4 58.5 24.2 71.7 48.3 19.2 40.8 Ours w/o mask 80.3 30.6 48.8 53.2 18.3 52.5 38.3 26.7 34.0 Ours w/o context 86.6 21.3 39.3 49.1 28.3 67.5 25.8 32.5 38.5
+Appendix G Simulation Evaluation Details
+G.1 SIMPLER Benchmark
 We evaluate on all eight SIMPLER [33] tasks in the visual matching setting, which include two robot platforms: Google Robot and WidowX.
@@
 In the case of GR00T [51], we use the official pretrained checkpoint and performe fine-tuning on the RT-1/Bridge dataset following the authors’ published guidelines accordingly.
-D.2 LIBERO
-The LIBERO benchmark [40] evaluates knowledge transfer in multitask and lifelong robot learning problems for robotic manipulation, consisting of four task suites: LIBERO-Spatial, LIBERO-Goal, LIBERO-Object, LIBERO-Long.
-For fine-tuning of our models, we reinitialize the linear state encoder, action encoder, and action decoder, and tune the full parameters (except for the vision encoder).
-We fine-tune all the models for 50k gradient steps on the fine-tuning dataset on each task suite of LIBERO.
-We follow the standard evaluation protocol to test by randomizing the initial states of the environments.
-Reported results (Table 3) are the average success rates across on a total of 200 trials and a total of 20 trials per task on each task suite.
-Please refer to Section 4.2 and Appendix C of LIBERO [40] for more details.
-Appendix E Real-world Robot Platforms Evaluation Details
-E.1 Realman robot arm
+Appendix H LIBERO Benchmark
+The LIBERO benchmark [39] evaluates knowledge transfer in multitask and lifelong robot learning problems for robotic manipulation, consisting of four task suites: LIBERO-Spatial evaluates the model’s performance under novel layouts with the same task and object types, LIBERO-Goal evaluates the model’s performance under novel tasks with the same object types and layouts, LIBERO-Object evaluates the model’s performance under novel object types with the same tasks and layouts, LIBERO-Long evaluates the model’s performance under diverse set of objects, layouts and backgrounds.
+Each task suite contains 10 tasks with 50 human demonstrations per task for fine-tuning.
+Baselines and Experimental Setup
+We compare with the following existing models: Diffusion Policy [11] trained from scratch, Octo [52], OpenVLA [29], $\pi_{0}$ [4], $\pi_{0}$ FAST [54], TraceVLA [71] and SpatialVLA  [55].
+For $\pi_{0}$, we use the open source version [57] and the same training set as our model.
+All models follow a two-stage pretraining-finetuning protocol.
+We finetune villa-X and villa-X w/o latent on the demonstration data of the each task suite separately, and test on the LIBERO simulator for 10 tasks and 20 trials per task on each task suite.
+Experimental Results
+Table 9 summarizes the success rates on each task suite of LIBERO.
+Our model achieves better performance than existing methods in all the four task suites.
+Also, our model with latent action achieves higher performance on all the four task suites and average performance, confirming that the proposed latent action expert improves the manipulation performance.
+Table 9: Evaluation on 4 LIBERO task suites of villa-X and existing methods.
+Method Spatial Object Goal Long Average Diffusion Policy [11] 78.3 92.5 68.3 50.5 72.4 Octo-base [52] 78.9 85.7 84.6 51.1 75.1 OpenVLA [29] 84.7 88.4 79.2 53.7 76.5 $\pi_{0}$ (reimplement [57]) 88.0 88.5 87.0 61.0 81.1 $\pi_{0}$-FAST [54] 96.4 96.8 88.6 60.2 85.5 TraceVLA [71] 84.6 85.2 75.1 54.1 74.8 SpatialVLA  [55] 88.2 89.9 78.6 55.5 78.1 Ours w/o latent 86.0 86.5 85.0 70.0 81.9 Ours 97.5 97.0 91.5 74.5 90.1
+Appendix I Real-world Robot Platforms Evaluation Details
+I.1 Realman robot arm
 The Realman robot arm setup is shown in Figure 5 (upper).
@@
 Figure 11: Realman evaluation trajectory examples.
-E.2 XHand dexterous hand
+I.2 XHand dexterous hand
 The Xhand setup is shown in Figure 5 (lower).